what does this problem mean? the larger of two numbers is 1 more than 3 imes the smaller.the difference between 8 times the smaller and 2 times the larger is 10
Answer by BurningFlame_marc(77)

Hello!

Actually, that question means that 3x+1=larger number and x=smaller number.
8x-2(3x+1)=10. that is what it means.
Now I will solve it.

8x-2(3x+1)=10
8x-6x-2=10
2x-2=10
2x=12
x=6


The smaller number is 6, the larger number is 19